# Securing Large Language Models: Threats, Vulnerabilities and Responsible Practices

### v1

@@ 1 Introduction
Recently, Large Language Models (LLMs) have initiated a significant paradigm shift in the areas of Natural Language Processing (NLP) including Natural Language Generation (NLG). LLMs are generally characterized by a large number of parameters -— usually ranging from millions to trillions – and are constructed using deep neural networks, mainly transformer architectures Vaswani et al. (2017); Lin et al. (2021). They undergo pre-training on massive amounts of text data, often collected from the web, and leverage self-supervised Jiao et al. (2023), semi-supervised Shi et al. (2023a), or Reinforcement Learning (RL) Ouyang et al. (2022); Gulcehre et al. (2023) methods for pre-training and fine-tuning. Researchers continue to explore ways to fine-tune and adapt these LLMs for specific applications, making them indispensable tools in the NLP community and beyond.
@@
To foster responsible and ethical use of LLMs, it is essential to develop methods and frameworks that can assess, improve, and govern LLMs in accordance with the principles of fairness, accountability, transparency, and explainability. This task requires a comprehensive and interdisciplinary investigation of LLMs from a security, ethical and risk mitigation perspective.
While there are several existing studies Ganguli et al. (2022); Huang et al. (2023b); Sun et al. (2023); Deshpande et al. (2023b); Wang et al. (2023a) that examine the security and risks related to LLMs, the fast-paced advancement and innovation in this domain calls for a more rigorous and systematic analysis.
With the goal of raising awareness and promoting responsible practices regarding LLMs, we explore the potential threats and vulnerabilities associated with LLMs and categorize them into model-based, training-time and inference-time vulnerabilities.
@@
  • Risk Mitigation Strategies: Our comprehensive assessment covers mitigation strategies like red teaming, model editing, watermarking, and AI-generated text detection techniques while discussing limitations and trade-offs.
@@ 2 Background
In order to enhance the clarity of this paper and avoid redundancy, we provide a glossary of the key terms that are frequently used in this work. We present table 1 that shows the terms, their concise definitions, and the paper sections where they are elaborated in depth. We encourage the readers to refer to the glossary whenever they encounter a term that is unfamiliar or unclear to them. The glossary is intended to serve as a quick reference guide and not as a comprehensive explanation of the concepts.
@@
A data pre-processing step for retraining or fine-tuning LLMs. It explicitly removes data points identified as vulnerable to leakage from dataset and retrains or fine-tunes LLMs on the processed dataset. Cao and Yang (2015)
3.1
Memorization
@@
Memorization is the phenomenon of LLMs retaining and reproducing information from their training data. Memorization can be beneficial for tasks that require factual or linguistic knowledge, but also problematic for privacy, security, and quality reasons Hartmann et al. (2023a); Zhong et al. (2023).
3.2
Association
@@
association is the ability of LLMs to form connections between different pieces of information, such as words, entities, concepts, or events. Association in LLMs can enable various applications, such as knowledge retrieval, text summarization, and question answering. However, association in LLMs can also pose challenges, such as privacy leakage, hallucination, and bias Shao et al. (2023); Du et al. (2023a); Chen and Ding (2023)
3.2
Auditing
@@
Auditing is to perform an examination to understand the implications and consequences of LLM memorization. For example, in auditing verbatim memorization, the examination would include a setup to generate arbitrary generated strings in order to detect if the LLM can provide an output of the said arbitrary generated string with high probability. Hartmann et al. (2023b)
3.2
Adversarial Attack
@@
An adversarial attack is a method that leverages the vulnerabilities or shortcomings of an LLM to induce erroneous or deceptive outputs. Adversarial attacks can be utilized for malicious purposes, such as creating misinformation, circumventing security protocols, or undermining the reliability of the model Shayegani et al. (2023).
4
Attack Success Rate (ASR)
@@
An attack success rate is a measure of the efficacy of an adversarial attack on a machine learning model. It is computed as the fraction of successful attacks over the total number of attacks. A successful attack is one that makes the model produce an erroneous prediction or output . An attack success rate can change depending on the type of attack, the type of model, the type of task, and the degree of perturbation Wu et al. (2021).
4
Model Extraction Attack
@@
A form of adversarial attack that leverages a large number of queries and their corresponding responses to extract the knowledge or parameters of an LLM. The extracted information can then be used to train a reduced parameter model that approximates the target LLM, or to conduct subsequent attacks on the LLM or other models. Prompt extraction Kirk et al. (2023), model leeching Birch et al. (2023a) and side channel attacks Tol and Sunar (2023) are common examples of model extraction attacks.
4.1.1
Data Poisoning
@@
Is an attack that corrupts the training data of an LLM, impacting its performance, behavior, or output. Data poisoning can lead to issues such as biases, falsehoods, toxicity, backdoors, or vulnerabilities in the model . Data poisoning can be deliberate by malicious actors who aim to harm or hijack the model, or accidental by negligent or uninformed data providers who neglect data quality and security standards. Data poisoning can be avoided or reduced by using reliable data sources, checking and cleaning the data, detecting anomalies in the model, and assessing the model for resilience Chen et al. (2017); Schwarzschild et al. (2020); Yang et al. (2021)
4.2.1
Backdoor Attack
@@
A type of malicious manipulation that embeds a hidden trigger in the model, causing it to perform normally on benign samples but exhibit degraded performance on poisoned ones. This issue is particularly concerning within communication networks where reliability and security are paramount Yang et al. (2023a). Input-triggered, instruction-triggered and demonstration-triggered are some common ways to launch a backdoor attack on LLMs Zhao et al. (2023a); Yao et al. (2023a); Huang et al. (2023a); Zhu et al. (2022).
4.2.2
Paraphrasing Attack
@@
An attack that uses a paraphraser model to rewrite AI-generated text and evade its detection. It can enhance the naturalness and human-likeness of the AI-generated text, and bypass the signatures or patterns of the detectors. A paraphrasing attack can challenge the security and reliability of LLMs and their applications Krishna et al. (2023); Sadasivan et al. (2023).
4.3.1
Spoofing Attack
@@
A spoofing attack in context of LLMs is an adversarial attack that imitates a specific LLM with an altered LLM to create similar outputs. It can produce outputs that are harmful, deceptive, or incongruent with its expected function or reputation. For instance, a spoofed chatbot can mimic popular LLMs and generate abusive and false utterances or disclose confidential information which endanger the security and privacy of LLM-based applications Shayegani et al. (2023).
4.3.1
Prompt Injection
@@
Is an adversarial attack that seeks to alter the output of an LLM by providing it with instructions that override or conflict with the intended ones. Liu et al. (2023a); Greshake et al. (2023)
4.3.2
Prompt Leaking
@@
Prompt leaking is a type of prompt injection, which is a malicious strategy that exploits the vulnerability of a language model to alter its output with deceptive prompts. Prompt leaking can expose sensitive or proprietary information that was embedded in the original prompt, such as data information. Prompt leaking can endanger the security and privacy of applications that rely on language models Perez and Ribeiro (2022).
4.3.2
Jailbreaking Attack
@@
A jailbreaking attack is a form of attack that exploits the vulnerability of a LLMs to alter its output with deceptive prompts. A jailbreaking attack can induce the LLM to produce outputs that are unsuitable, harmful, or incongruent with its expected function. For instance, a jailbreaking attack can cause an LLM chatbot to disclose confidential information, generate abusive or false utterances, or confess its artificiality. Jailbreaking attacks can jeopardize the security and privacy of LLM-based applications Zhang et al. (2023b); Deng et al. (2023b).
4.3.3
Black-box Detection
@@
Black-box detection is the task of identifying inaccuracies in the outputs of LLMs or detecting LLM-generated texts without accessing their internal states or training data. It typically involves asking follow-up questions, analysing the model’s responses, and applying a classifier to detect patterns of deception Anonymous (2023). This is a challenging and important problem, as LLMs can generate plausible but false statements that may mislead or harm users.
6.3
White-box Detection
@@
The task of detecting LLM-generated texts by having full access to the target model. This method can prevent unauthorized use of LLMs and monitor their generation behavior Wang et al. (2023f)
6.3
Watermarking
@@
Watermarking in LLMs is a technique that embeds hidden signals in the text generated by an LLM to make it algorithmically identifiable as synthetic, while being imperceptible to humans. Watermarking can help mitigate the potential risks of LLMs, such as disinformation, plagiarism, or impersonation, by proving the ownership, authenticity, and integrity of the text Kirchenbauer et al. (2023a); Tang et al. (2023a)
6.3.4
@@
LLM Security Study Security & Privacy Risks 3 Adversarial Risks 4 Misuse Risks 5 Risk Mitigation Strategies 6 Leaking Sensitive Information 3.1 Memorizing Training Data 3.2 Security Holes in LLM-Generated Codes 3.3 Model-based Vulnerabilities 4.1 Inference-Time Vulnerabilities 4.3 Training-Time Vulnerabilities 4.2 Factual Inconsistency & Unreliability 5.1 Generating Toxic & Discriminatory Contents 5.2 Copyright Infringements 5.3 Plagiarisms 5.3 Spreading Misinformation 5.4 Editing LLMs 6.1 Cybersecurity Resilience Engineering 6.2 Detecting LLM Generated Text 6.3 Model Extraction Attacks 4.1.1 Model Leeching Attacks 4.1.1 Model Imitation Attacks 4.1.2 Data Poisoning Attacks 4.2.1 Backdoor Attacks 4.2.2 Paraphrasing Attacks 4.3.1 Spoofing Attacks 4.3.1 Prompt Injection Attacks 4.3.2 Jailbreaking Attacks 4.3.3 Post-training Gradient Editing 6.1 Post-Training Weight Editing 6.1 Memory-Based Model Editing 6.1 Ensemble of Model Editing 6.1 Red Teaming 6.2 Green Teaming 6.2 Supervised Methods 6.3.1 Zero-Shot Methods 6.3.2 Retrieval-Based Methods 6.3.3 Watermarking Methods 6.3.4 Feature-Based Methods 6.3.5
@@ 3 Security and Privacy Concerns of LLMs
LLMs are powerful tools that may pose several security risks for enterprises and individuals. In this section, we explore critical concerns, including sensitive information leakage, memorizing training data and potential security holes in generated code. By understanding these risks, we can foster responsible AI usage and enhance overall security practices.
@@ 3.1 Sensitive Information Leakage
Data leakage occurs when sensitive, personal, or private information from the training data or real user input is exposed through the model’s completion. For example, if a language model generates a credit card number or an email address that belongs to a real identity, it is considered data leakage.
On the other hand, memorization refers to the ability of a language model to recall specific examples from the training data and reproduce them during inference. For instance, if the model generates the exact text of an article or a news headline it was trained on, data memorization has occurred. It’s important to note that data leakage can be caused by memorization, but this is not always the case.
@@ 3.2 Memorizing Training Data
The emergence of LLMs, coupled with their extensive parameter count, has raised concerns about the extent of their memorization. Specifically, when these models learn from training data, there is a question of whether the information contained in that data can be accessed through prompts that request access to the inner mechanisms of the LLM.
@@
In conclusion, as LLMs are becoming larger in terms of number of parameters, it is crucial to comprehend their memorization. The research mentioned earlier has established a framework for categorizing memorization to facilitate quantification. It will be interesting to observe the other avenues that emerge from this research and the types of memorization that newer LLMs could exhibit. Table 2 shows different memorization studies along with datasets and techniques that are used for memorization.
@@
Hartmann et al. Hartmann et al. (2023b), for example, introduce the concept of “Auditing” to discern whether an LLM merely engages in verbatim memorization or leverages that information for deeper insights. The authors further present a comprehensive taxonomy that categorizes different forms of memorization within LLMs, including verbatim text, facts, ideals, algorithms, writing styles, distributional properties, and alignment goals.
@@
Memorization pertains to the model’s capacity to memorize sensitive data and subsequently retrieve it in response to user queries. Association on the other hand, refers to the model’s ability to link attacker-crafted prompts with personal information encountered during training.
To quantify memorization, the Huang et al. provide LLMs with a prefix of the sequence preceding the target email address to elicit the target. To quantify association, they designed different prompts based on how email addresses appear in a sentence.
Their findings reveales that while LLMs did indeed leak private data due to memorization, their performance in association tasks was comparatively weak. Interestingly, the risk escalated with model size, aligning with expectations—larger models exhibited enhanced sophistication and memory capabilities.

### v2

@@ 1 Introduction
Recently, Large Language Models (LLMs) have initiated a significant paradigm shift in the areas of Natural Language Processing (NLP) including Natural Language Generation (NLG). LLMs are generally characterized by a large number of parameters – usually ranging from millions to trillions – and are constructed using deep neural networks, mainly transformer architectures Vaswani et al. (2017); Lin et al. (2021). They undergo pre-training on massive amounts of text data, often collected from the web, and leverage self-supervised Jiao et al. (2023), semi-supervised Shi et al. (2023a), or Reinforcement Learning (RL) Ouyang et al. (2022); Gulcehre et al. (2023) methods for pre-training and fine-tuning. Researchers continue to explore ways to fine-tune and adapt these LLMs for specific applications, making them indispensable tools in the NLP community and beyond.
@@
To foster responsible and ethical use of LLMs, it is essential to develop methods and frameworks that assess, improve, and govern LLMs in accordance with the principles of fairness, accountability, transparency, and explainability. This requires a comprehensive and interdisciplinary investigation of LLMs from a security, ethical and risk mitigation perspective.
While there are several existing studies Ganguli et al. (2022); Huang et al. (2023b); Sun et al. (2023); Deshpande et al. (2023b); Wang et al. (2023a) that examine the security and risks of LLMs, the fast-paced advancement and innovation in this domain calls for a rigorous and systematic analysis.
With the goal of raising awareness and promoting responsible practices, we explore the potential threats and vulnerabilities associated with LLMs and categorize them into model-based, training-time and inference-time vulnerabilities.
@@
  • Risk Mitigation Strategies: Our comprehensive assessment covers mitigation strategies like red teaming, model editing, watermarking, and AI-generated text detection techniques including discussions on limitations and trade-offs.
@@ 2 Background
In order to enhance the clarity of this paper and avoid redundancy, we provide a glossary of the key terms that are frequently used in this work. We present table LABEL:tab:glossary that shows the terms, their concise definitions, and the paper sections where they are elaborated in depth. We encourage the readers to refer to the glossary whenever they encounter a term that is unfamiliar or unclear to them. The glossary is intended to serve as a quick reference guide and not as a comprehensive explanation of the concepts.
@@
Unlearning A data pre-processing step for retraining or fine-tuning LLMs. It explicitly removes data points identified as vulnerable to leakage from dataset and retrains or fine-tunes LLMs on the processed dataset. Cao and Yang (2015); Chakraborty et al. (2024) 3.1
@@
Memorization Memorization is the phenomenon of LLMs retaining and reproducing information from their training data. Memorization can be beneficial for tasks that require factual or linguistic knowledge, but also problematic for privacy, security, and quality reasons Hartmann et al. (2023a); Zhong et al. (2023). 3.2
@@
Association association is the ability of LLMs to form connections between different pieces of information, such as words, entities, concepts, or events. Association in LLMs can enable various applications, such as knowledge retrieval, text summarization, and question answering. However, association in LLMs can also pose challenges, such as privacy leakage, hallucination, and bias Shao et al. (2023); Du et al. (2023a); Chen and Ding (2023) 3.2
@@
Auditing Auditing is to perform an examination to understand the implications and consequences of LLM memorization. For example, in auditing verbatim memorization, the examination would include a setup to generate arbitrary generated strings in order to detect if the LLM can provide an output of the said arbitrary generated string with high probability. Hartmann et al. (2023b) 3.2
Emergent Misalignment refers to unintended, unpredictable behavior that arises during fine-tuning, leading models to exhibit unsafe or unaligned responses across multiple dimensions. Betley et al. (2025) 3.3
@@
Adversarial Attack An adversarial attack is a method that leverages the vulnerabilities or shortcomings of an LLM to induce erroneous or deceptive outputs. Adversarial attacks can be utilized for malicious purposes, such as creating misinformation, circumventing security protocols, or undermining the reliability of the model Bachu et al. (2024); Shayegani et al. (2023). 4
@@
Attack Success Rate (ASR) An attack success rate is a measure of the efficacy of an adversarial attack on a machine learning model. It is computed as the fraction of successful attacks over the total number of attacks. A successful attack is one that makes the model produce an erroneous prediction or output . An attack success rate can change depending on the type of attack, the type of model, the type of task, and the degree of perturbation Wu et al. (2021). 4
@@
Model Extraction Attack A form of adversarial attack that leverages a large number of queries and their corresponding responses to extract the knowledge or parameters of an LLM. The extracted information can then be used to train a reduced parameter model that approximates the target LLM, or to conduct subsequent attacks on the LLM or other models. Prompt extraction Kirk et al. (2023), model leeching Birch et al. (2023a) and side channel attacks Tol and Sunar (2023) are common examples of model extraction attacks. 4.1.1
@@
Data Poisoning Is an attack that corrupts the training data of an LLM, impacting its performance, behavior, or output. Data poisoning can lead to issues such as biases, falsehoods, toxicity, backdoors, or vulnerabilities in the model . Data poisoning can be deliberate by malicious actors who aim to harm or hijack the model, or accidental by negligent or uninformed data providers who neglect data quality and security standards. Data poisoning can be avoided or reduced by using reliable data sources, checking and cleaning the data, detecting anomalies in the model, and assessing the model for resilience Chen et al. (2017); Schwarzschild et al. (2020); Yang et al. (2021) 4.2.1
@@
Backdoor Attack A type of malicious manipulation that embeds a hidden trigger in the model, causing it to perform normally on benign samples but exhibit degraded performance on poisoned ones. This issue is particularly concerning within communication networks where reliability and security are paramount Yang et al. (2023a). Input-triggered, instruction-triggered and demonstration-triggered are some common ways to launch a backdoor attack on LLMs Zhao et al. (2023a); Yao et al. (2023a); Huang et al. (2023a); Zhu et al. (2022). 4.2.2
@@
Paraphrasing Attack An attack that uses a paraphraser model to rewrite AI-generated text and evade its detection. It can enhance the naturalness and human-likeness of the AI-generated text, and bypass the signatures or patterns of the detectors. A paraphrasing attack can challenge the security and reliability of LLMs and their applications Shayegani et al. (2025); Krishna et al. (2023); Sadasivan et al. (2023). 4.3.1
@@
Spoofing Attack A spoofing attack in context of LLMs is an adversarial attack that imitates a specific LLM with an altered LLM to create similar outputs. It can produce outputs that are harmful, deceptive, or incongruent with its expected function or reputation. For instance, a spoofed chatbot can mimic popular LLMs and generate abusive and false utterances or disclose confidential information which endanger the security and privacy of LLM-based applications Shayegani et al. (2023). 4.3.1
@@
Prompt Injection Is an adversarial attack that seeks to alter the output of an LLM by providing it with instructions that override or conflict with the intended ones. Liu et al. (2023a); Greshake et al. (2023) 4.3.2
@@
Prompt Leaking Prompt leaking is a type of prompt injection, which is a malicious strategy that exploits the vulnerability of a language model to alter its output with deceptive prompts. Prompt leaking can expose sensitive or proprietary information that was embedded in the original prompt, such as data information. Prompt leaking can endanger the security and privacy of applications that rely on language models Perez and Ribeiro (2022). 4.3.2
@@
Jailbreaking Attack A jailbreaking attack is a form of attack that exploits the vulnerability of a LLMs to alter its output with deceptive prompts. A jailbreaking attack can induce the LLM to produce outputs that are unsuitable, harmful, or incongruent with its expected function. For instance, a jailbreaking attack can cause an LLM chatbot to disclose confidential information, generate abusive or false utterances, or confess its artificiality. Jailbreaking attacks can jeopardize the security and privacy of LLM-based applications Shayegani et al. (2024); Zhang et al. (2023b); Deng et al. (2023b). 4.3.3
@@
Black-box Detection Black-box detection is the task of identifying inaccuracies in the outputs of LLMs or detecting LLM-generated texts without accessing their internal states or training data. It typically involves asking follow-up questions, analysing the model’s responses, and applying a classifier to detect patterns of deception Anonymous (2023). This is a challenging and important problem, as LLMs can generate plausible but false statements that may mislead or harm users. 6.3
@@
White-box Detection The task of detecting LLM-generated texts by having full access to the target model. This method can prevent unauthorized use of LLMs and monitor their generation behavior Wang et al. (2023f) 6.3
@@
Watermarking Watermarking in LLMs is a technique that embeds hidden signals in the text generated by an LLM to make it algorithmically identifiable as synthetic, while being imperceptible to humans. Watermarking can help mitigate the potential risks of LLMs, such as disinformation, plagiarism, or impersonation, by proving the ownership, authenticity, and integrity of the text Kirchenbauer et al. (2023a); Tang et al. (2023a) 6.3.4
@@
LLM Security Study Security & Privacy Risks 3 Adversarial Risks 4 Misuse Risks 5 Risk Mitigation Strategies 6 Leaking Sensitive Information 3.1 Memorizing Training Data 3.2 Security Holes in LLM-Generated Codes 3.3 Model-based Vulnerabilities 4.1 Inference-Time Vulnerabilities 4.3 Training-Time Vulnerabilities 4.2 Factual Inconsistency & Unreliability 5.1 Generating Toxic & Discriminatory Contents 5.2 Copyright Infringements 5.3 Plagiarisms 5.3 Spreading Misinformation 5.4 Editing LLMs 6.1 Cybersecurity Resilience Engineering 6.2 Detecting LLM Generated Text 6.3 Model Extraction Attacks 4.1.1 Model Leeching Attacks 4.1.1 Model Imitation Attacks 4.1.2 Data Poisoning Attacks 4.2.1 Backdoor Attacks 4.2.2 Paraphrasing Attacks 4.3.1 Spoofing Attacks 4.3.1 Prompt Injection Attacks 4.3.2 Jailbreaking Attacks 4.3.3 Post-training Gradient Editing 6.1 Post-Training Weight Editing 6.1 Memory-Based Model Editing 6.1 Ensemble of Model Editing 6.1 Red Teaming 6.2 Blue Teaming 6.2 Green Teaming 6.2 Purple Teaming 6.2 Supervised Methods 6.3.1 Zero-Shot Methods 6.3.2 Retrieval-Based Methods 6.3.3 Watermarking Methods 6.3.4 Feature-Based Methods 6.3.5
@@ 3 Security and Privacy Concerns of LLMs
LLMs are powerful tools, but may pose security risks for both enterprises and individuals. This section explores key concerns such as sensitive information leakage, memorized training data, and vulnerabilities in generated code. Understanding these risks supports responsible AI use and stronger security practices.
@@ 3.1 Sensitive Information Leakage
Data leakage occurs when sensitive, personal, or private information from training data or real user input is exposed through the model’s completion. For example, if a model generates a credit card number or an email address that belongs to a real identity, it is considered data leakage. Figure 2 illustrates an example of data leakage.
Conversely, memorization refers to a language model’s tendency to recall and reproduce specific examples from its training data during inference. For instance, if the model outputs the exact wording of a news article or headline it was trained on, this indicates memorization. Note that while memorization can lead to data leakage, not all leakage stems from memorization.
@@ 3.2 Memorizing Training Data
The rise of LLMs and their vast parameter counts has raised concerns about how much they memorize. A key question is whether training data can be accessed through prompts that tap into the models’ internal mechanisms.
@@
Figure 3: Similarities and differences between memorization, association and data leakage.
@@
In conclusion, as LLMs are becoming larger in terms of number of parameters, it is crucial to comprehend their memorization. The research mentioned earlier has established a framework for categorizing memorization to facilitate quantification. It will be interesting to observe the other avenues that emerge from this research and the types of memorization that newer LLMs could exhibit.
@@
Hartmann et al. Hartmann et al. (2023b), for example, introduce the concept of “Auditing” to discern whether an LLM merely engages in verbatim memorization or leverages that information for deeper insights. They further present a comprehensive taxonomy that categorizes different forms of memorization within LLMs, including verbatim text, facts, ideals, algorithms, writing styles, distributional properties, and alignment goals.
@@
Memorization pertains to the model’s capacity to memorize sensitive data and subsequently retrieve it in response to user queries. Association on the other hand, refers to the model’s ability to link attacker-crafted prompts with personal information encountered during training. Figure 3 demonstrates the similarities and differences between these concepts.
To quantify memorization, Huang et al. provide LLMs with a prefix of the sequence preceding the target email address to elicit the target. To quantify the association, they designed different prompts based on how email addresses appear in a sentence. Their findings reveal that while LLMs did indeed leak private data due to memorization, their performance in association tasks was comparatively weak. Interestingly, risk escalated with model size, aligning with expectations: larger models exhibited enhanced sophistication and memory capabilities.
Recent work such as Stoehr et al. (2024) have expanded on the verbatim memorization for prefix-lengths and have went into memorization for token paragraphs. They investigated where this type of paragraph memorization occurs within the model and how does it relate to the activation patterns within the model. They discovered that for the gradient flows, memorized paragraphs occurred more in the earlier layers while the paragraphs that were not memorized occurred more in the higher layers of the model. Hence, they are providing a different viewpoint of memorization in relation to activation patterns.
@@
Mitigation Strategies The ability for LLMs to memorize the training data has been a great concern and there have been new and novel approaches to prevent it. Kassem et al. Kassem et al. (2023) utilize an RL paraphrasing policy to demonstrate that it can reduce memorization. Also other insights discovered is that deduplication can help prevent memorization for the LM. From a different viewpoint, Ozadayi et al. Ozdayi et al. (2023) leverage prompt tuning in order to assess the ability of extracting training data memorization from LLMs. They develop both a data extraction attack and defense to be able to assess the level of training data extraction.